OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 +bx+c=0. -sprt(7), 4sprt(7)

OpenStudy (anonymous):

\[(x+\sqrt{7})(x-4\sqrt{7})=0\]

OpenStudy (anonymous):

multiply out get \[x^2-3 \sqrt{7} x-28=0\]